# Propaganda to Hate: A Multimodal Analysis of Arabic Memes with Multi-Agent LLMs

### v1

@@ Abstract
In the past decade, social media platforms have been used for information dissemination and consumption. While a major portion of the content is posted to promote citizen journalism and public awareness, some content is posted to mislead users. Among different content types such as text, images, and videos, memes (text overlaid on images) are particularly prevalent and can serve as powerful vehicles for propaganda, hate, and humor. In the current literature, there have been efforts to individually detect such content in memes. However, the study of their intersection is very limited. In this study, we explore the intersection between propaganda and hate in memes using a multi-agent LLM-based approach. We extend the propagandistic meme dataset with coarse and fine-grained hate labels. Our finding suggests that there is an association between propaganda and hate in memes. We provide detailed experimental results that can serve as a baseline for future studies. We will make the experimental resources publicly available to the community.
@@ 1 Introduction
Our research is situated at the intersection of several rapidly evolving fields: multimodal content analysis, propaganda detection, and hate speech identification. While significant strides have been made in these areas for English and other high-resource languages, the Arabic digital landscape remains comparatively understudied. Our work extends the efforts of [5] and [18] in Arabic propaganda detection, incorporating their insights into our analysis of meme content. By focusing on Arabic memes, we address the gap in the literature The outcomes of this research can benefit social media platforms, policymakers, and civil society organizations working to combat the spread of harmful content online.
The key aspects of our work are as follows: (i) we present a novel multi-agent LLM-based approach to analyze the association between propaganda and hateful memes in Arabic social media content; (ii) we demonstrate the application of multi-agent LLM systems for automated annotation of complex, multimodal data, offering a scalable solution for processing large volumes of data in low-resource settings; (iii) in addition to coarse-grained hateful categories, we also explore fine-grained categorization of hateful and non-hateful memes; (iv) we provide experimental results on both coarse and fine-grained categories; (v) finally, we will make the labeled dataset of hateful memes available to the community.
@@ 2.1 Propagandistic Content
Textual Content: The study of propagandistic content has attracted significant attention in recent years. Da et al. (2020) introduced a large-scale dataset for fine-grained propaganda detection in news articles, presenting a corpus of 350K sentences annotated with 18 propaganda techniques [8]. The annotation schema has been extended to include 23 techniques and a multilingual corpus development has been proposed in [27]. Following the same annotation schema, datasets and shared tasks have been developed for Arabic [4, 20, 19].
Multimodal Content: Building on previous research in textual content, Dimitrov et al. (2021) introduced SemEval-2021 Task 6, which focuses on detecting persuasion techniques in texts and images within memes [14]. Subsequently, the focus has expanded to include the detection of multilingual and multimodal propagandistic memes [12]. Similarly, related work on Arabic involves the development of datasets and a shared task for propaganda detection [3, 21]. Fang et al. (2022) utilized separate networks to embed text and images and then fused these multi-modal embeddings. They employed a split-and-share module with multi-level representations to enhance the detection performance of persuasive techniques [15].
@@ 2.2 Hateful Memes
The study of hateful memes presents unique challenges due to their multimodal nature. Kiela et al. (2020) introduced the Hateful Memes Challenge, a large-scale dataset and benchmark for multimodal hate detection. This work highlighted the importance of integrating both textual and visual elements to identify hate speech in memes [23]. Sharma et al. (2022) proposed a multi-task learning approach for detecting hate speech in memes by combining visual and textual features. Their model demonstrated improved performance over unimodal approaches, underscoring the synergy between different modalities in hate speech detection [30]. Addressing the challenge of low-resource languages, datasets have also been developed in various languages [22]. In [31], the authors provide a detailed survey of multimodal and harmful memes, highlighting the significance of the problem and proposing future research avenues.
@@ 3.1 Propagandistic Memes
Figure 1: Experimental pipeline in which LLM agents function as annotators and a consolidator.
@@ 3.2 Hatefulness and Fine-grained Categories
We used ArAIEval-2024 dataset [21] for the hatefulness and their fine-grained categorization. Our motivation to use ArAIEval-2024 dataset is that this is the only meme dataset currently available for Arabic, which has already been annotated for propagandistic content. Another motivation was to understand the association between propagandistic and hateful memes. In Figure 1, we provide full pipeline for the data preparation to classification experiments.
@@ 3.2.1 LLM Agents as Annotators
To employ LLM agents as annotators, we selected three well-known and top-performing commercial models: OpenAI’s GPT-4o [26], Google’s Gemini Pro (version 1.5) [32], and Claude 3.5 (Sonnet). [^2] https://www.anthropic.com/news/claude-3-5-sonnet.
For the annotation process, we use the same manual procedure discussed in [18], which involves a two-phase approach. In the first phase, known as the annotation phase, three annotators independently annotate memes following the guidelines outlined in 3.2.2. In the second phase, known as the consolidation phase, we review and resolve any disagreements from the annotations received during the first phase. As illustrated in the figure, highlighted in dark red, we employ LLM agents as annotators in the first phase and as a consolidator in the second phase. For each phase, we use a specific prompt in a zero-shot setup for the LLM agent. Following the annotation guidelines discussed below, we ask an LLM agent to perform two tasks: (Task 1) label each meme as hateful or not-hateful, and (Task 2) based on the label from Task 1, provide a fine-grained categorization. For example, if a meme is categorized as hateful in Task 1, it should then provide a fine-grained label from one of the eight categories mentioned below. The prompt in the second phase is slightly different. Here, the task also involves considering the labels obtained from the first phase to make a final decision. For this phase, we have experimented with using GPT-4o as the consolidator.
@@ 3.2.3 Annotation Instructions
The purpose of this annotation is to identify whether a meme is hateful or not-hateful. A hateful meme can attack different individuals, organizations or entities. Therefore, another task is identifying the attack types. A non-hateful meme can be humorous or sarcastic. Therefore, the idea to also identify the sub-categories within non-hateful memes. Below we provide the definitions:
@@
A direct or indirect attack on people based on characteristics, including ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, and disability or disease. We define attack as violent or dehumanizing (comparing people to non-human things, e.g., animals) speech, statements of inferiority, and calls for exclusion or segregation. Mocking hate crime is also considered hate speech. Attacking groups perpetuating hate (e.g. terrorist groups) is not considered hate.
@@
  • Dehumanizing: Explicitly or implicitly describing or presenting a group as subhuman.
  • Inferiority: Claiming that a group is inferior, less worthy or less important than either society in general or another group
  • Inciting violence: Explicitly or implicitly calling for harm to be inflicted on a group, including physical attacks
  • Mocking: Making jokes about, undermining, belittling, or disparaging a group
  • Contempt: Expressing intensely negative feelings or emotions about a group
  • Slurs: Using prejudicial terms to refer to, describe or characterize a group
  • Exclusion: Advocating, planning or justifying the exclusion or segregation of a group from all of society or certain parts
  • Other: None of the above
@@ 4 Experiments
Text classification: We extracted text from propagandistic memes and applied various text classification techniques such as AraBERT, mBERT, CAMelBERT, and Qarib-BERT ([6, 11, 1]). The original dataset is imbalanced, and so we implemented a class weighting scheme during the fine-tuning process. Moreover, we optimized the model by adjusting the dropout rate. This approach led to significant improvements over using the original dataset alone. We embedded the LoRa to fine-tune the model in an efficient way that does not require fine-tuning all the parameters of the model. However, embedding LoRA in the final classification layer did not improve the performance. We then fine-tuned AraBERT model.
@@
Multimodal classification: To extract visual and text features, we applied ConvNext tiny and AraBERT models respectively. We combine the features using a fusion layer. We froze the visual models and trained the classification layer with textual data. We also adjusted the dropout rate to improve the performance.
Experimental Setup: We performed all of our experiments and trained our models on an Nvidia-RTX 2080 GPU. We employed the Adam optimizer with an initial learning rate of 1e-5 and 1e-4 for text and image, respectively. We used a batch size of 32 containing a 128-character sentence length and 224x224 image. We set the dropout rate to 0.25 for text data and trained for 50 epochs. For image data, we set a dropout rate of 0.5 and trained for 30 epochs. For multimodal data, we trained the model for 100 epochs with a stochastic drop rate of 0.2. Note that our choice of the models and parameters was inspired by prior studies [3, 29].
5 Results and Discussion
In Table 3, we present the results for different classification setups. For hate vs. not-hate classification across different modalities, considering macro-F1, the text and multimodal models exhibit similar performance. For the multimodal model, we obtained model with a macro-F1 score of 0.709 and an accuracy of 0.764. For the text modality, we obtained a macro-F1 score of 0.705 and an accuracy of 0.818. For the image modality, we obtained a macro-F1 score of 0.669 and an accuracy of 0.775.
To understand the class imbalance issue, we selected 500 propaganda labels and 500 non-propaganda labels from the dataset and applied the fusion model. This yielded a macro-F1 score of 0.709 and an accuracy of 0.818. The hateful memes were further fine-grained into eight labels, while the non-hateful memes were fine-grained into two labels, as shown in Table 2. We applied the fusion model individually to the hateful and non-hateful fine-grained labels. The F1-score for the hateful fine-grained memes is 0.224, with an accuracy of 0.166, which is very low. This occurs because there are multiple labels within the hateful category, and the dataset is imbalanced for fine-grained hateful memes. In contrast, with only two labels in the non-hateful memes, the model performs better in classification, achieving a macro-F1 score of 0.537 and an accuracy of 0.622.

### v2

@@ Abstract
In the past decade, social media platforms have been used for information dissemination and consumption. While a major portion of the content is posted to promote citizen journalism and public awareness, some content is posted to mislead users. Among different content types such as text, images, and videos, memes (text overlaid on images) are particularly prevalent and can serve as powerful vehicles for propaganda, hate, and humor. In the current literature, there have been efforts to individually detect such content in memes. However, the study of their intersection is very limited. In this study, we explore the intersection between propaganda and hate in memes using a multi-agent LLM-based approach. We extend the propagandistic meme dataset with coarse and fine-grained hate labels. Our finding suggests that there is an association between propaganda and hate in memes. We provide detailed experimental results that can serve as a baseline for future studies. We will make the experimental resources publicly available to the community. [^1] https://github.com/firojalam/propaganda-and-hateful-memes.git
@@ 1 Introduction
Our research lies at the intersection of multimodal content analysis, propaganda detection, and hate speech identification. While progress has been made in these fields for English and other high-resource languages, the research for Arabic remains underexplored. Building on the work of [5] and [18] in Arabic propaganda detection, our study analyzes Arabic memes, addressing a gap in the literature. The findings can assist social media platforms, policymakers, and civil society organizations in combating harmful online content.
The key aspects of our work are as follows: (i) we present a novel multi-agent LLM-based approach to analyze the association between propaganda and hateful memes in Arabic social media content; (ii) we demonstrate the application of multi-agent LLM systems for automated annotation of complex, multimodal data, offering a scalable solution for processing large volumes of data in low-resource settings; (iii) in addition to coarse-grained hateful categories, we also explore fine-grained categorization of hateful and non-hateful memes; (iv) we provide experimental results on both coarse and fine-grained categories; (v) finally, we will make the dataset of hateful memes available to the community.
@@ 2.1 Propagandistic Content
Textual Content: The study of propagandistic content has attracted significant attention in recent years. Da et al. (2020) introduced a large-scale dataset for fine-grained propaganda detection in news articles, presenting a corpus of 350K sentences annotated with 18 propaganda techniques [8]. The annotation schema has been extended to include 23 techniques and a multilingual corpus has been proposed in [28]. Following the same annotation schema, datasets have been developed for Arabic and shared tasks has been organized [4, 20, 19].
Multimodal Content: Building on previous research in textual content, Dimitrov et al. (2021) introduced SemEval-2021 Task 6, which focuses on detecting persuasion techniques in texts and images within memes [14]. Subsequently, the focus has expanded to include the detection of multilingual and multimodal propagandistic memes [12]. Similarly, related work on Arabic involves the development of datasets and a shared task for propaganda detection [3, 21]. Fang et al. (2022) used separate networks to embed text and images, fusing these multi-modal embeddings. A split-and-share module with multi-level representations was employed to improve persuasive technique detection [15].
@@ 2.2 Hateful Memes
The study of hateful memes presents unique challenges due to their multimodal nature. Kiela et al. (2020) introduced the Hateful Memes Challenge, a large-scale dataset and benchmark for multimodal hate detection. This work highlighted the importance of integrating both textual and visual elements to identify hate speech in memes [23]. Addressing the challenge of low-resource languages, datasets have also been developed in various languages [22]. In [31], the authors provide a detailed survey of multimodal and harmful memes, highlighting the significance of the problem and proposing future research avenues.
@@ 3.1 Propagandistic Memes
Figure 1: Experimental pipeline with LLM agents as annotators and consolidator.
@@ 3.2 Hatefulness and Fine-grained Categories
For the hatefulness and their fine-grained categorization we used ArAIEval-2024 dataset, mentioned earlier. Our motivation to use ArAIEval-2024 dataset is that this is the only meme dataset currently available for Arabic, which has already been annotated for propagandistic content. Another motivation was to understand the association between propagandistic and hateful memes. In Figure 1, we provide full pipeline for the data preparation to classification experiments.
@@ 3.2.1 LLM Agents as Annotators
To employ LLM agents as annotators, we selected three well-known and top-performing commercial models: OpenAI’s GPT-4o [27], Google’s Gemini Pro (version 1.5) [32], and Claude 3.5 (Sonnet). [^3] https://www.anthropic.com/news/claude-3-5-sonnet. For the annotation process, we use the same manual procedure discussed in [18], which involves a two-phase approach. In the first phase, known as the annotation phase, three annotators independently annotate memes following the guidelines outlined in 3.2.2. In the second phase, known as the consolidation phase, we review and resolve any disagreements from the annotations received during the first phase. As illustrated in the figure, highlighted in dark red, we employ LLM agents as annotators in the first phase and as a consolidator in the second phase. For each phase, we use a specific prompt in a zero-shot setup for the LLM agent. Following the annotation guidelines discussed below, we ask an LLM agent to perform two tasks: (Task 1) label each meme as hateful or not-hateful, and (Task 2) based on the label from Task 1, provide a fine-grained categorization. For example, if a meme is categorized as hateful in Task 1, it should then provide a fine-grained label from one of the eight categories mentioned below. The prompt in the second phase is slightly different. Here, the task also involves considering the labels obtained from the first phase to make a final decision. For this phase, we have experimented with using GPT-4o as the consolidator.
@@ 3.2.3 Annotation Instructions
The purpose of this annotation is to identify whether a meme is hateful or not-hateful. A hateful meme can attack different individuals, organizations or entities. Therefore, another task is identifying the attack types. A non-hateful meme can be humorous or sarcastic. Therefore, the idea to also identify the sub-categories within non-hateful memes. We adopted the annotation definition and instructions from prior work [23, 25]. Below we provide the definitions:
@@
A direct or indirect attack on individuals based on characteristics such as ethnicity, race, nationality, immigration status, religion, caste, sex, gender identity, sexual orientation, disability, or disease. We define an attack as hate-speech that is violent, dehumanizing (such as comparing individuals to non-human entities like animals), involves statements of inferiority, or calls for exclusion or segregation. Mocking hate crimes is also classified as hate speech. However, attacks directed at groups that perpetuate hate (e.g., terrorist organizations) are not considered hate speech. [^4] https://transparency.meta.com/en-gb/policies/community-standards/hate-speech/
@@
  • Dehumanizing: Explicitly or implicitly portraying or describing a group as subhuman.
  • Inferiority: Asserting that a group is inferior, less worthy, or less important than society as a whole or compared to another group.
  • Inciting violence: Explicitly or implicitly advocating for harm to be inflicted on a group, including physical violence.
  • Mocking: Joking about, ridiculing, demeaning, or disparaging a group.
  • Contempt: Expressing strong negative emotions or feelings toward a group.
  • Slurs: Using biased or derogatory terms to refer to, describe, or characterize a group.
  • Exclusion: Advocating for, planning, or justifying the exclusion or segregation of a group from society as a whole or from specific areas.
  • Other: None of the above.
@@ 4 Experiments
Text classification: We extracted text from propagandistic memes and applied various text classification techniques such as AraBERT, mBERT, CAMelBERT, and Qarib-BERT ([6, 11, 1]). The original dataset is imbalanced, and so we implemented a class weighting scheme during the fine-tuning process. Moreover, we optimized the model by adjusting the dropout rate. This approach led to significant improvements over using the original dataset alone. We embedded the LoRa to fine-tune the model in an efficient way that does not require fine-tuning all the parameters of the model. However, embedding LoRA did not improve the performance. We then fine-tuned AraBERT model.
@@
Multimodal classification: To extract visual and text features, we applied ConvNext tiny and AraBERT models respectively. We combine the features using a fusion layer. We froze the visual models and trained the classification layer with textual data. We also used a dropout rate to improve the performance.
Experimental Setup: We performed all of our experiments and trained our models on an Nvidia-RTX 2080 GPU. We employed the Adam optimizer with an initial learning rate of 1e-5 and 1e-4 for text and image, respectively. We used a batch size of 32, a sequence length of 128. We set the dropout rate to 0.25 for text data and trained for 50 epochs. For image data, we set a dropout rate of 0.5 and trained for 30 epochs. For multimodal data, we trained the model for 100 epochs with a stochastic drop rate of 0.2. Note that our choice of the models and parameters was inspired by prior studies [3, 30].
@@
5 Results and Discussion
In Table 3, we present the results for different classification setups. For hate vs. not-hate classification across different modalities, considering macro-F1, the text and multimodal models exhibit similar performance. For the multimodal model, we obtained model with a macro-F1 of 0.709 and an accuracy of 0.764. For the text modality, we obtained a macro-F1 of 0.705 and an accuracy of 0.818. For the image modality, we obtained a macro-F1 of 0.669 and an accuracy of 0.775.
To understand the class imbalance issue, we selected 500 propaganda labels and 500 non-propaganda labels from the dataset and applied the fusion model. This yielded a macro-F1 of 0.709 and an accuracy of 0.818. The hateful memes were further fine-grained into eight labels, while the non-hateful memes were fine-grained into two labels, as shown in Table 2. We applied the fusion model individually to the hateful and non-hateful fine-grained labels. The F1-score for the hateful fine-grained memes is 0.224, with an accuracy of 0.166, which is very low. This occurs because there are multiple labels within the hateful category, and the dataset is imbalanced for fine-grained hateful memes. In contrast, with only two labels in the non-hateful memes, the model performs better in classification, achieving a macro-F1 of 0.537 and an accuracy of 0.622.